# From Formal Language Theory to Statistical Learning: Finite Observability of Subregular Languages

### v1

@@ Abstract
We prove that all standard subregular language classes are linearly separable when represented by their deciding predicates. This establishes finite observability and guarantees learnability with simple linear models. Synthetic experiments confirm perfect separability under noise-free conditions, while real-data experiments on English morphology show that learned features align with well-known linguistic constraints. These results demonstrate that the subregular hierarchy provides a rigorous and interpretable foundation for modeling natural language structure. Our code used in real-data experiments is available at https://github.com/UTokyo-HayashiLab/subregular.
@@ 1 Introduction
In recent decades, research has increasingly focused on the subregular hierarchy, a set of language classes strictly contained within the regular languages but capable of capturing a wide range of phonological, morphological, and even syntactic phenomena (Heinz, 2010; Rogers et al., 2013; Heinz et al., 2011; Graf, 2017; Graf and Mayer, 2018; Graf, 2022b, a; Torres et al., 2023; Hanson, 2023). This line of work has revealed that linguistic constraints often fall within these low-complexity classes, explaining their learnability and robustness in acquisition (Heinz and Idsardi, 2011; Saffran et al., 1996). For example, phonotactic restrictions such as vowel harmony, consonant dissimilation, and tonal patterns have been successfully modeled with subregular languages (Rogers et al., 2013; Heinz, 2010; Jardine and Heinz, 2016). Morphological and syntactic dependencies have likewise been shown to align with subregular complexity (Graf, 2017; Chandlee, 2014).
While the algebraic and logical characterizations of subregular classes are well established (Straubing, 1994; Thomas, 1997; Pin, 1995), there has been limited work providing a geometric perspective on why these classes matter for learnability. At the same time, statistical learning theory has established powerful results connecting linear separability to generalization guarantees via the VC dimension, perceptron mistake bounds, and margin-based analyses (Vapnik, 1998; Blumer et al., 1989; Novikoff, 1962; Cortes and Vapnik, 1995; Shalev-Shwartz and Ben-David, 2014). These results suggest that if subregular classes can be shown to be linearly separable in appropriate finite feature spaces, then we can directly apply statistical learning theory to natural language phenomena.
@@
    We prove that all standard subregular classes (SL, SP, LT, PT, LTT, TSL) are finitely observable (see Appendix A.4).
@@ 2 Related Work
The subregular hierarchy has been established as a family of language classes that are both computationally weaker than regular languages and yet linguistically expressive enough to capture many phonological and morphological constraints. The algebraic characterization of star-free languages and their correspondence with first-order logic FO[<<] (via the classical results of McNaughton, Papert, and Schützenberger) laid much of the theoretical foundation (McNaughton and Papert, 1971; Straubing, 1994; Schützenberger, 1965; Eilenberg, 1974; Pin, 1995; Thomas, 1997). In particular, Simon’s characterization of Piecewise Testable (PT) languages (Simon, 1975) and the algebraic and logical perspectives on Locally Testable (LT) languages (McNaughton and Papert, 1971; Straubing, 1994; Pin, 1995; Thomas, 1997) illustrate how language membership can be determined by finite local observations, thereby connecting naturally to SLkk and SPkk. Rogers et al. (2013) systematized the subregular complexity of phonological patterns, while Heinz (2010) showed how long-distance phonotactic restrictions are learnable precisely because they fall within the subregular hierarchy. Tier-based Strictly Local (TSL) languages, which localize nonlocal dependencies via tier projection, have proven to be a powerful extension (Heinz et al., 2011), with applications extending to syntax as well (Graf, 2017; Graf and Mayer, 2018). Our contribution complements this tradition by providing a geometric characterization of subregular classes via linear separability, offering an alternative perspective to classical algebraic and automata-theoretic accounts.
Linguistic evidence and applications (phonology, morphology, syntax).
Subregular classes have been applied to model local and nonlocal constraints across phonology, morphology, and syntax. In phonology, Strictly Local languages capture cluster restrictions and syllable-structure constraints (Rogers et al., 2013), while Strictly Piecewise languages capture nonadjacent harmony processes (Heinz, 2010). TSL languages handle long-distance vowel harmony and other tier-based phenomena (Heinz et al., 2011; Jardine and Heinz, 2016), while tonal and assimilation/dissimilation processes have also been argued to fall within subregular subclasses (Graf and Mayer, 2018; Avcu et al., 2017). In morphology, affix ordering restrictions and stem–affix co-occurrence patterns have been modeled using SL, LT, and TSL languages (Rogers et al., 2013; Chandlee, 2014). In syntax, dependencies have been reanalyzed in terms of Strictly Local and Strictly Piecewise constraints (Graf, 2017). The present work connects these descriptive successes to interpretable linear models, providing a theoretical justification for lightweight learning and inference modules grounded in subregular structure.
@@
The grammatical inference literature has long studied the identifiability of language classes, beginning with Gold’s identification-in-the-limit framework (Gold, 1967), Angluin’s query learning paradigm (Angluin, 1987), and consolidated in de la Higuera’s monograph (de la Higuera, 2010). Early work showed that certain subclasses of regular languages (e.g., $k$-testable in the strict sense) admit efficient inference algorithms (García and Vidal, 1990; Oncina and García, 1992). For subregular languages, numerous learnability results have been established for SL, SP, and TSL classes (Rogers et al., 2013; Heinz, 2010; Heinz et al., 2011; Graf and Mayer, 2018; Jardine and Heinz, 2016; Chandlee, 2014). Meanwhile, statistical learning theory provides a different angle: linear separability yields strong guarantees via the VC dimension (Vapnik, 1998; Blumer et al., 1989; Shalev-Shwartz and Ben-David, 2014), the perceptron mistake bound (Novikoff, 1962), and SVM margin analyses (Cortes and Vapnik, 1995). Our contribution demonstrates that subregular classes are linearly separable in appropriate finite feature spaces, thereby allowing these guarantees to apply directly to natural language classes themselves. This unifies classical grammatical inference with modern statistical learning theory.
Connections between logic, algebra, and geometry.
The logical and algebraic characterizations of regular languages (FO[<<], monoids, and aperiodicity (McNaughton and Papert, 1971; Straubing, 1994; Schützenberger, 1965; Pin, 1995; Thomas, 1997; Bojańczyk, 2014)) have traditionally been studied in symbolic and automata-theoretic terms. Recent work has refined the complexity and decidability results for locally testable languages, piecewise testable languages, and their relatives (Place and Zeitoun, 2014; Place et al., 2013). Our contribution provides a complementary geometric perspective, recasting these classes as halfspaces in the Boolean hypercube via minterm embeddings. Specifically, any class definable by a finite set of primitive predicates (substrings, subsequences, prefixes, thresholds) is linearly separable with constant margin, and the standard subregular families fall neatly into this framework; a bridge from algebra and logic to statistical learning and convex geometry.
@@ 3.1 Predicates and Finite Observability
  • Substring predicates (SLkk): for $g\in(\Sigma\cup\{\#\})^{k}$, $p_{g}(x)=\mathbf{1}[\,g\text{ occurs contiguously in }\tilde{x}\,$]. Example: in an SL3 language forbidding 3-gram ngt, the constraint is captured by requiring $p_{\texttt{ngt}}(x)=0$.
  • Subsequence predicates (SPkk): for $h\in(\Sigma\cup\{\#\})^{k}$, $q_{h}(x)=\mathbf{1}[\,h\text{ occurs as a subsequence of }\tilde{x}\,$]. Example: in SP2, $q_{\texttt{a\_i}}(x)$ indicates whether a precedes i; forbidding such pairs encodes vowel-harmony style constraints.
  • Boundary predicates (LTkk): for $u,v\in(\Sigma\cup\{\#\})^{k-1}$, $\pi_{u}(x)=\mathbf{1}[\,\mathrm{prefix}_{k-1}(\tilde{x})=u\,]$, $\sigma_{v}(x)=\mathbf{1}[\,\mathrm{suffix}_{k-1}(\tilde{x})=v\,]$.
[…]
@@ 3.3 Main Theorem: Finite Observability $\Rightarrow$ Linear Separability
Each feature vector $\Phi(x)$ is one-hot with $\|\Phi(x)\|_{2}=1$, so the effective radius is $R=1$ and the geometric margin is $\gamma\geq 1/(2\|w\|_{2})$. Hence a perceptron run on separable data suffers at most $M\leq(1/\gamma)^{2}\leq 4\|w\|_{2}^{2}=4|S|$ mistakes (Novikoff, 1962). Moreover, the class of halfspaces in $\mathbb{R}^{2^{n}}$ has VC dimension at most $2^{n}$ (Vapnik, 1998; Blumer et al., 1989), and margin-based generalization bounds apply directly (Cortes and Vapnik, 1995; Shalev-Shwartz and Ben-David, 2014). In short, finitely observable classes are easy to learn under standard linear methods.
@@ 3.4 From Theory to Practice: Low-dimensional Predicate Features
  • SLkk: presence/absence of $k$-grams (with boundary padding) suffices to implement forbidden-substring constraints.
  • SPkk: presence/absence of $k$-length subsequences suffices to implement forbidden nonadjacent pairs/tuples.
  • LTkk: $k$-gram features augmented with (k1)(k{-}1)-length prefix/suffix indicators determine membership.
  • PTmm: presence/absence of subsequences of length $\leq m$ (Simon’s theorem (Simon, 1975)).
[…]
  • TSLkk: $k$-gram features on tier projections $\pi_{T}$ (for $T\subseteq\Sigma$) capture nonlocal constraints through local ones on the tier (Heinz et al., 2011).
@@
Our results require fixing class parameters (e.g., $k$, $m$, $\boldsymbol{\tau}$) and the finite alphabet, which is standard when defining target language families. Unions over unbounded parameters (e.g., PT $\bigcup_{m}\mathrm{PT}_{m}$) should be viewed as directed unions of finitely observable subclasses (Appendix A.4). While the minterm embedding itself is exponential, practical predicate sets for the standard subregular families are polynomial in $|\Sigma|$ and the fixed parameters, and can be trained with sparse linear methods.
@@ 4 A Counterexample Showing the Limits of Fixed Finite Predicate Sets
Our main results hinge on finite observability: once a finite set of predicates $P=\{p_{1},\dots,p_{n}\}$ is fixed, any language definable in this framework is a union of the finitely many predicate cells (truth assignments) in $\{0,1\}^{n}$. This section shows that, for any fixed finite predicate set $P$, there exist regular languages that cannot be detected (i.e., represented as unions of predicate cells), hence cannot be linearly separated in the corresponding feature space. The proof relies on the Myhill–Nerode theorem (Sipser, 2012).
@@
It is well known (and follows from Myhill–Nerode) that $L_{m}$ has exactly $m$ distinct right-congruence classes, corresponding to length modulo $m$ (Sipser, 2012). For any equivalence relation $E$ whose index (number of classes) is strictly smaller than $m$, $E$ cannot refine the Myhill–Nerode congruence of $L_{m}$; hence $L_{m}$ cannot be a union of $E$-classes. Since $\equiv_{P}$ has at most $2^{n}$ classes and $m>2^{n}$, $\equiv_{P}$ does not refine the Myhill–Nerode congruence of $L_{m}$. Therefore $L_{m}$ is not a union of $\equiv_{P}$-classes and cannot be represented in the predicate framework determined by $P$. Because linear classification over $P$ can only realize unions of $\equiv_{P}$-cells, no linear separator exists for $L_{m}$ in that fixed feature space. ∎
@@
Theorem 4.1 formalizes an intuitive limitation: fixing a finite observation vocabulary $P$ yields at most $2^{n}$ observable profiles, while regular languages can require arbitrarily many distinct right-congruence classes (e.g., $m$ for $L_{m}$), even over a unary alphabet. Thus, unlike the subregular families discussed in this paper, the class of all regular languages cannot be captured (nor linearly separated) within any single finite predicate space. This clarifies the scope of our finite-observability program: it targets families whose essential structure is visible through a bounded set of primitive observations.
Star-Free Languages as an Intermediate Case.
In addition to regular languages, it is instructive to consider star-free languages. They are strictly weaker than the full class of regular languages but strictly more expressive than the standard subregular families. Classical results show that locally testable and piecewise testable languages are contained within star-free, while strictly local and strictly piecewise families form even smaller subclasses McNaughton and Papert (1971); Brzozowski and Simon (1973); Simon (1975); Rogers et al. (2007). Thus we have a proper hierarchy:
$\text{SL},\ \text{SP}\;\subset\;\text{LT},\ \text{PT}\;\subset\;\text{Star-Free}\;\subset\;\text{Regular}.$
Unlike the subregular classes, the star-free family does not admit finite observability. This follows from the fact that FO[<], which characterizes star-free languages, forms strict quantifier-rank and dot-depth hierarchies Brzozowski and Cohen (1971); Kufleitner and Lauser (2012); Place and Zeitoun (2021). No finite predicate basis suffices to capture all star-free languages. Thus, star-free languages provide a useful counterexample: finite observability holds robustly for the subregular families but breaks down once we move upward in expressive power, even before reaching all regular languages.
@@ 5 Synthetic Experiments
A) SL3 (noise) B) SP2 (noise) C) LTT2 (noise)
@@
D) SL3 (train size) E) SP2 (train size) F) LTT2 (train size)
@@
Figure 1: Accuracy of synthetic classification: (A) SL3, (B) SP2 and (C) LTT2 under label noise, (D) SL3, (E) SP2 and (F) LTT2 with varying training sizes.
@@ 5.1 Linear Separability
To empirically validate our theoretical results, we conducted controlled experiments on artificially generated languages. The goal of these experiments is not to maximize accuracy per se, but rather to visualize how linear separability manifests across different subregular classes when noise or sample size is varied.
To keep the presentation concise and focused, we concentrate on three representative classes from the subregular hierarchy: Strictly Local (SL), Strictly Piecewise (SP), and Locally Threshold Testable (LTT). SL captures local adjacency constraints and provides the most transparent connection to phonotactic well-formedness. SP models long-distance restrictions, such as co-occurrence bans, which are common in phonology and morphology. Finally, LTT extends local tests with counting thresholds, allowing us to probe more complex phenomena and to examine how separability behaves as the constraints become richer.
@@
5.1.1 Strictly Local Languages (SLk)

### v2

@@ Abstract
We prove that all standard subregular language classes are linearly separable when represented by their deciding predicates. This establishes finite observability and guarantees learnability with simple linear models. Synthetic experiments confirm perfect separability under noise-free conditions, while real-data experiments on English morphology show that learned features align with well-known linguistic constraints. These results demonstrate that the subregular hierarchy provides a rigorous and interpretable foundation for modeling natural language structure. Our code used in real-data experiments is available at https://github.com/{Anonymous}.
@@ 1 Introduction
In recent decades, research has increasingly focused on the subregular hierarchy, a set of language classes strictly contained within the regular languages but capable of capturing a wide range of phonological, morphological, and even syntactic phenomena (Heinz, 2010; Heinz et al., 2011; Graf, 2022b, a; Torres et al., 2023; Hanson, 2024). This line of work has revealed that linguistic constraints often fall within these low-complexity classes, explaining their learnability and robustness in acquisition (Heinz and Idsardi, 2011; Saffran et al., 1996). For example, phonotactic restrictions such as vowel harmony, consonant dissimilation, and tonal patterns have been successfully modeled with subregular languages (Heinz, 2010; Jardine and Heinz, 2016). Morphological and syntactic dependencies have likewise been shown to align with subregular complexity (Chandlee, 2014).
While the algebraic and logical characterizations of subregular classes are well established (Straubing, 2012; Thomas, 1997; Pin, 1996), there has been limited work providing a geometric perspective on why these classes matter for learnability. At the same time, statistical learning theory has established powerful results connecting linear separability to generalization guarantees via the VC dimension, perceptron mistake bounds, and margin-based analyses (Vapnik, 1999; Blumer et al., 1989; Novikoff, 1962; Cortes and Vapnik, 1995; Shalev-Shwartz and Ben-David, 2014). These results suggest that if subregular classes can be shown to be linearly separable in appropriate finite feature spaces, then we can directly apply statistical learning theory to natural language phenomena.
@@
    We prove that all standard subregular classes ($\mathsf{SL}$, $\mathsf{SP}$, $\mathsf{LT}$, $\mathsf{PT}$, $\mathsf{LTT}$, $\mathsf{TSL}$) are finitely observable (see Appendix A.4).
@@ 2 Related Work
The subregular hierarchy has been established as a family of language classes that are both computationally weaker than regular languages and yet linguistically expressive enough to capture many phonological and morphological constraints. The algebraic characterization of star-free languages and their correspondence with first-order logic FO[<<] laid much of the theoretical foundation (McNaughton and Papert, 1971; Straubing, 2012; Schützenberger, 1965; Eilenberg, 1974; Thomas, 1997). In particular, Simon’s characterization of Piecewise Testable $\mathsf{(PT)}$ languages (Simon, 1975) and the algebraic and logical perspectives on Locally Testable $\mathsf{(LT)}$ languages (McNaughton and Papert, 1971; Straubing, 2012; Thomas, 1997) illustrate how language membership can be determined by finite local observations, thereby connecting naturally to $\mathsf{SL}_{k}$ and $\mathsf{SP}_{k}$. Tier-based Strictly Local ($\mathsf{TSL}$) languages, which localize nonlocal dependencies via tier projection, have proven to be a powerful extension (Heinz et al., 2011). Our contribution complements this tradition by providing a geometric characterization of subregular classes via linear separability, offering an alternative perspective to classical algebraic and automata-theoretic accounts.
@@
The grammatical inference literature has long studied the identifiability of language classes, beginning with Gold’s identification-in-the-limit framework (Gold, 1967), Angluin’s query learning paradigm (Angluin, 1987), and consolidated in de la Higuera’s monograph (de la Higuera, 2010). Early work showed that certain subclasses of regular languages (e.g., $k$-testable in the strict sense) admit efficient inference algorithms (García and Vidal, 1990; Oncina and García, 1992). For subregular languages, numerous learnability results have been established for $\mathsf{SL}$, $\mathsf{SP}$, and $\mathsf{TSL}$ classes (Heinz, 2010; Heinz et al., 2011; Jardine and Heinz, 2016; Chandlee, 2014). Meanwhile, statistical learning theory provides a different angle: linear separability yields strong guarantees via the VC dimension (Vapnik, 1999; Blumer et al., 1989; Shalev-Shwartz and Ben-David, 2014), the perceptron mistake bound (Novikoff, 1962), and SVM margin analyses (Cortes and Vapnik, 1995). Our contribution demonstrates that subregular classes are linearly separable in appropriate finite feature spaces, thereby allowing these guarantees to apply directly to natural language classes themselves. This unifies classical grammatical inference with modern statistical learning theory.
@@ 3.1 Predicates and Finite Observability
  • Substring predicates ($\mathsf{SL}_{k}$): for $g\in(\Sigma\cup\{\#\})^{k}$, $p_{g}(x)=\mathbf{1}[\,g\text{ occurs contiguously in }\tilde{x}\,$]. Example: in an SL3 language forbidding 3-gram ngt, the constraint is captured by requiring $p_{\texttt{ngt}}(x)=0$.
  • Subsequence predicates ($\mathsf{SP}_{k}$): for $h\in(\Sigma\cup\{\#\})^{k}$, $q_{h}(x)=\mathbf{1}[\,h\text{ occurs as a subsequence of }\tilde{x}\,$]. Example: in SP2, $q_{\texttt{a\_i}}(x)$ indicates whether a precedes i; forbidding such pairs encodes vowel-harmony style constraints.
  • Boundary predicates ($\mathsf{LT}_{k}$): for $u,v\in(\Sigma\cup\{\#\})^{k-1}$, $\pi_{u}(x)=\mathbf{1}[\,\mathrm{prefix}_{k-1}(\tilde{x})=u\,]$, $\sigma_{v}(x)=\mathbf{1}[\,\mathrm{suffix}_{k-1}(\tilde{x})=v\,]$.
  • Thresholded count predicates ($\mathsf{LTT}_{k,\boldsymbol{\tau}}$): for $g$ with $|g|\leq k$ and $t\in\{1,\dots,\tau_{g}\}$, $c_{g,t}(x)=\mathbf{1}[\,\#\{g\text{ in }\tilde{x}\}\geq t\,]$.
@@ 3.3 Main Theorem: Finite Observability $\Rightarrow$ Linear Separability
Each feature vector $\Phi(x)$ is one-hot with $\|\Phi(x)\|_{2}=1$, so the effective radius is $R=1$ and the geometric margin is $\gamma\geq 1/(2\|w\|_{2})$. Hence a perceptron run on separable data suffers at most $M\leq(1/\gamma)^{2}\leq 4\|w\|_{2}^{2}=4|S|$ mistakes (Novikoff, 1962). Moreover, the class of halfspaces in $\mathbb{R}^{2^{n}}$ has VC dimension at most $2^{n}$ (Blumer et al., 1989), and margin-based generalization bounds apply directly (Cortes and Vapnik, 1995; Shalev-Shwartz and Ben-David, 2014). In short, finitely observable classes are easy to learn under standard linear methods.
@@ 3.4 From Theory to Practice: Low-dimensional Predicate Features
  • $\mathsf{SL}_{k}$: presence/absence of $k$-grams (with boundary padding) suffices to implement forbidden-substring constraints.
  • $\mathsf{SP}_{k}$: presence/absence of $k$-length subsequences suffices to implement forbidden nonadjacent pairs/tuples.
  • $\mathsf{LT}_{k}$: $k$-gram features augmented with (k1)(k{-}1)-length prefix/suffix indicators determine membership.
  • $\mathsf{PT}_{m}$: presence/absence of subsequences of length $\leq m$ (Simon, 1975).
  • $\mathsf{LTT}_{k,\boldsymbol{\tau}}$: bounded-threshold counts for substrings of length $\leq k$.
  • $\mathsf{TSL}_{k}$: $k$-gram features on tier projections $\pi_{T}$ (for $T\subseteq\Sigma$) capture nonlocal constraints through local ones on the tier (Heinz et al., 2011).
@@
Our results require fixing class parameters (e.g., $k$, $m$, $\boldsymbol{\tau}$) and the finite alphabet, which is standard when defining target language families. Unions over unbounded parameters (e.g., $\mathsf{PT}$ $\bigcup_{m}\mathsf{PT}_{m}$) should be viewed as directed unions of finitely observable subclasses (Appendix A.4). While the minterm embedding itself is exponential, practical predicate sets for the standard subregular families are polynomial in $|\Sigma|$ and the fixed parameters, and can be trained with sparse linear methods.
@@ 4 A Counterexample Showing the Limits of Fixed Finite Predicate Sets
Our main results hinge on finite observability: once a finite set of predicates $P=\{p_{1},\dots,p_{n}\}$ is fixed, any language definable in this framework is a union of the finitely many predicate cells (truth assignments) in $\{0,1\}^{n}$. This section shows that, for any fixed finite predicate set $P$, there exist regular languages that cannot be detected (i.e., represented as unions of predicate cells), hence cannot be linearly separated in the corresponding feature space. The proof relies on the Myhill–Nerode theorem (Sipser, 1996).
@@
It is well known (and follows from Myhill–Nerode) that $L_{m}$ has exactly $m$ distinct right-congruence classes, corresponding to length modulo $m$ (Sipser, 1996). For any equivalence relation $E$ whose index (number of classes) is strictly smaller than $m$, $E$ cannot refine the Myhill–Nerode congruence of $L_{m}$; hence $L_{m}$ cannot be a union of $E$-classes. Since $\equiv_{P}$ has at most $2^{n}$ classes and $m>2^{n}$, $\equiv_{P}$ does not refine the Myhill–Nerode congruence of $L_{m}$. Therefore $L_{m}$ is not a union of $\equiv_{P}$-classes and cannot be represented in the predicate framework determined by $P$. Because linear classification over $P$ can only realize unions of $\equiv_{P}$-cells, no linear separator exists for $L_{m}$ in that fixed feature space. ∎
@@
Theorem 4.1 formalizes an intuitive limitation: fixing a finite observation vocabulary $P$ yields at most $2^{n}$ observable profiles, while regular languages can require arbitrarily many distinct right-congruence classes (e.g., $m$ for $L_{m}$), even over a unary alphabet. Thus, unlike the subregular families discussed in this paper, the class of all regular languages cannot be captured (nor linearly separated) within any single finite predicate space.
@@ 5 Synthetic Experiments
A) $\mathsf{SL}_{3}$ (noise) B) $\mathsf{SP}_{2}$ (noise) C) $\mathsf{LTT}_{2}$ (noise)
@@
D) $\mathsf{SL}_{3}$ (train size) E) $\mathsf{SP}_{2}$ (train size) F) $\mathsf{LTT}_{2}$ (train size)
@@
Figure 1: Accuracy of synthetic classification: (A) $\mathsf{SL}_{3}$, (B) $\mathsf{SP}_{2}$ and (C) $\mathsf{LTT}_{2}$ under label noise, (D) $\mathsf{SL}_{3}$, (E) $\mathsf{SP}_{2}$ and (F) $\mathsf{LTT}_{2}$ with varying training sizes.
@@ 5.1 Linear Separability
To empirically validate our theoretical results, we conducted controlled experiments on artificially generated languages. The goal of these experiments is not to maximize accuracy itself, but rather to visualize how linear separability manifests across different subregular classes when noise or sample size is varied.
To keep the presentation concise and focused, we concentrate on three representative classes from the subregular hierarchy: Strictly Local ($\mathsf{SL}$), Strictly Piecewise ($\mathsf{SP}$), and Locally Threshold Testable ($\mathsf{LTT}$). $\mathsf{SL}$ captures local adjacency constraints and provides the most transparent connection to phonotactic well-formedness. $\mathsf{SP}$ models long-distance restrictions, such as co-occurrence bans, which are common in phonology and morphology. Finally, $\mathsf{LTT}$ extends local tests with counting thresholds, allowing us to probe more complex phenomena and to examine how separability behaves as the constraints become richer.
@@
5.1.1 Strictly Local Languages ($\mathsf{SL}_{k}$)